If a function is defined as f(x,n) mod n. Determine the

i.    Domain of f

ii.   Range of f

iii.         G(g(g(g(7)))) if g (n) = f(209, n).


1
Expert's answer
2021-09-21T12:35:06-0400

i)

Domain: "x\isin Z"


ii)

Range: "f(x,n)\isin Z"


iii)

"g(7)=f(209,7)=6"

"g(g(7))=g(6)=f(209,6)=5"

"g(g(g(7)))=g(5)=f(209,5)=4"

"G(g(g(g(7))))=g(4)=f(209,4)=1"